# Canonical formalism in the Schwarzschild spacetime

### v2

@@ II.3.3 Summary
where the notation $y_{\mathcal{N}}^{\textrm{A}}$ includes an A for “Andoyer”. They are in a 1-to-1 correspondence with the old canonical set $y_{\mathcal{N}}$ defined in Sec. I.2.3 via Eqs. (59) and (60), and they generalize the classical Schwarzschild geodesic elements (cf. Sec. II.1).
@@ II.4.1 Andoyer Hamiltonian
Despite the striking similarity with the first integrals of the geodesic case (cf. the discussion below Eq. (52)), we would like to emphasize two points. First, all these first integrals include a spin contribution, and are not equal to their geodesic counterpart (despite their name). Second, the constant plane, orthogonal to $\vec{L}$, is not the orbital plane, in contrast to the geodesic case. Generic spinning orbits are not planar, but they do stay $O(\epsilon)$ to the plane orthogonal to $\vec{L}$, which can be considered to be the “mean” orbital plane, cf. Fig. 3.
@@ II.4.2 Constraints and equations of motion on $\mathcal{N}$
where we made the dependence on the Rüdiger invariant $\mathfrak{K}$ (cf. Eq. (44a)) explicit, even though, on $\mathcal{N}$, this quantity is merely function of the Andoyer variables; it is neither a coordinate nor a first integral. However, it reveals the practical advantage of choosing $C^{0}$ and $C^{1}$ over other choices for the TD SSC constraints: $C^{0}$ and $C^{1}$ can be directly related with the Rüdiger invariant $\mathfrak{K}$, and will be very useful on $\mathcal{P}$.
@@ III.1 Hamiltonian constraints
Two points must be made regarding the right-hand side of the aforementioned formula in Eq. (72). First, the non-vanishing 161616 The formula in Eq. (72) is singular when $\{C^{0},C^{1}\}=0$, which occurs when $\pi_{\alpha}=0$ [cf. Eq. (79)]. The singularity is only apparent, however, because both $\{C^{0},C^{1}\}$ and the numerator of the formula is always proportional to $\pi_{\alpha}$, which we will see from Eq. (186). The limit $\pi_{\alpha}\rightarrow 0$ of Eq. (72) is therefore regular. of the denominator $\{C_{0},C_{1}\}$ is a consequence of the symplectic nature of the Poisson structure on $\mathcal{P}$, as explained in Sec. V.A.2 of Paper I. Second, the right-hand side must be first computed with the brackets on $\mathcal{N}$, and then simplified the result, making use of the two constraints $C_{0}=0$ and $C_{1}=0$ if necessary. It is important to appreciate the order of these (non-commuting) operations; the redundant variables in Eq. (72) must be removed only after the brackets on $\mathcal{N}$ are completely evaluated.
@@ III.1.1 Andoyer spin variables
$\Delta^{2}$ $:=\sqrt{1-S_{\circ}^{2}/\pi_{\zeta}^{2}}\,\bigl{(}{\pi_{s}}\sin\zeta\cos s+{\pi_{s}}\cos\zeta\sin s\bigr{)}\,,$ (73e)
@@
We note that these Andoyer spin variables $(\vec{\Sigma},\vec{\Delta})$ are identical to $(\vec{S},\vec{D})$ in Eq. (28), except that the pair $(\sigma,\pi_{\sigma})$ is replaced with $(s,\pi_{s})$. Since $s$ is a combination (cf. Eq. (59)) of the orbital angles $\theta,\phi$ and the spin angle $\sigma$ adapted to Schwarzschild’s spherical symmetry, $(\vec{\Sigma},\vec{\Delta})$ will be much more practical to use than $(\vec{S},\vec{D})$, as we shall see.
@@ III.1.2 Choice of 10 variables on $\mathcal{P}$
Our next step is to chose $10$ variables to cover the phase space $\mathcal{P}$. Any subset of the $12$ Andoyer variables $y_{\mathcal{N}}^{\text{A}}$ (or any combination thereof) on $\mathcal{N}$ is in principle possible, altough in practice there is no general rule as to which variables should be retained or discarded. Our approach will be to retain the constants of motion $(E,\,L_{z},\,L^{2}=\mathfrak{Q},\,\mathfrak{K})$ (or, again, any combination thereof) introduced in Sec. I.3. Indeed, since we know that the system is integrable on $\mathcal{P}$ (cf. Sec. VI in Paper I), these constants of motion can be readily promoted to momenta on $\mathcal{P}$. We just have to express them in terms of Andoyer variables. These expressions can readily be found below Eq. (66) for $(E,L_{z},L)$, while that for $\mathfrak{K}$ is simply an upgrade from Eq. (44a) to
@@ III.1.4 $\mathcal{P}$-brackets
We can now compute the $\mathcal{P}$-brackets of all pairs of the $10$ variables of $y_{\mathcal{P}}^{\text{C}}$ in Eq. (77) via Eq. (72). To compute the $\mathcal{N}$-brackets in the right-hand side in the formula, we follow the general strategy adopted in the calculation of $\{C^{0},C^{1}\}$ in the preceding section; first, apply the Leibniz rule to relate the expression to the brackets in terms of $(\Sigma^{I},\Delta^{I})$, discarding any terms that are quadratic or higher in spins. Next, employ the SO$(1,3),\mathcal{N}$-brackets in Eq. (75) to eliminate the brackets in the expression. Finally, replace $(\Sigma^{I},\Delta^{I})$ with $\Gamma$ and $\pi_{\varpi}$ (using the relation in Eq. (184) of Appendix D.1), to put the result in the form solely expressed in terms of $y_{\mathcal{P}}^{\text{C}}$ in Eq. (77). The key technical point here is to fully make use of the Andyer spin variables $(\Sigma^{I},\Delta^{I})$ in the intermediate stage, which we find to be the most optimal. [^17] In principle, it is possible to directly calculate the $\mathcal{P}$-brackets from the Andoyer variables in Eq. (65), without relying on $(\Sigma^{I},\Delta^{I})$. However, it appears that such a strategy is computationally intractable in general (even with a computer algebra system).
An explicit listing of the $\mathcal{N}$-brackets between the constraints $C^{0}$ and $C^{1}$ in Eq. (78) and the $10$ variables $y_{\mathcal{P}}^{\text{C}}$ in Eq. (77) is given in Appendix D.1. Collecting the results from Eq. (79) and Eq. (186), and then inserting them into Eq. (72), we arrive at
@@ III.1.5 Generalized law of motion
where we used the Leibniz rule at the second equality. It is then straightforward to explicitly evaluate this law of motion, inserting the $\mathcal{P}$ brackets in Eq. (85) within Eq. (87). Performing the computation, we obtain181818 As expected, we also get the trivial equations $\{\mathcal{I},H_{\mathcal{P}}\}^{\mathcal{P}}=0$ for all first integrals $\mathcal{I}\in(\nu,\pi_{t},\pi_{\ell},\pi_{\nu},\pi_{\varpi})$.
@@ III.2.4 Fixing the residual dependence
The determination of $\tilde{t}_{3},\tilde{\pi}_{3},\tilde{\ell}_{3}$ and ${\varpi}_{3}$ as a function of variables $(r,\pi_{t},\pi_{\ell},\pi_{\varpi})$ is instead based on an alternative approach, which we detail in Appendix. D.2. The main conclusion of this analysis is that there is no unique solution, due to the freedom in choosing canonical coordinates in symplectic mechanics. Nevertheless, there is a trivial homogeneous solution expressed as (cf Eqs. (189) and (190))
@@
$y_{\mathcal{P}}^{\text{F}}=(T,r,\nu,\tilde{\ell},{\varpi},\pi_{t},\tilde{\pi}_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi})\,,$ (109)
@@ IV.1 The canonical Hamiltonian system
The first five of them carry the same physical interpretations as those in the geodesic case, and only $S_{\parallel}$ appears as a new integral of motion specific to the spinning case. In this sense, $H_{c}$ is a natural generalization of the well-established (covariant) geodesic Hamiltonian introduced in, e.g., Refs. [46, 53] [cf. Eq. (159) in Appendix. B]. It is important to note that only five out of the six integrals of motion are in pairwise involution, [^19] Two first integrals $\mathcal{I}_{1},\mathcal{I}_{2}$ are in involution whenever $\{\mathcal{I}_{1},\mathcal{I}_{2}\}=0$. because the canonical pair $(\nu,\,\pi_{\nu})$ satisfies bracket $\{\nu,\pi_{\nu}\}=1$ by definition.
@@ IV.2.1 Explicit formulae for the TD spin vector
IV.2.2 Comparison with the Marck representation
This subsection contains some results from the (so named) Marck formalism [34] — the parallel transport of the TD spins along the background geodesic orbit $\mathscr{L}_{0}$, formulated beyond the realm of the Hamiltonian mechanics. A short introduction to the Marck formalism is given in Appendix. C.
The meaning of the spin canonical variables $({\varpi},\,\pi_{\varpi})$ when viewed in Schwarzschild spacetime can be further clarified by the comparison between $S_{a}^{\text{\tiny{TD}}}(\tilde{x}^{i},\tilde{\pi}_{i})$ and an alternative representation of the TD spins $S_{a}^{\text{\tiny{TD}}}$ in terms of the Marck tetrad [cf. Eqs. (178) and (179)] which is parallel transported along the (background) Schwarzschild geodesics $\mathscr{L}_{0}$. It is helpful to note that the Marck tetrad is distinct from the natural-basis tetrad introduced in Eq. (33) and utilized in our Hamiltonian formulation. Also, the difference between the true orbit $\mathscr{L}$ and $\mathscr{L}_{0}$ is negligible here because $S_{a}^{\text{\tiny{TD}}}$ is already a quantity of $O(\epsilon)$.
An explicit form of $S_{a}^{\text{\tiny{TD}}}$ in the Marck tetrad is provided in Eq. (183), and the relevant expressions to the comparison with Eq. (126) and (127) are
$\pi_{\varpi}={\mathcal{S}}_{\parallel}\quad\text{with}\quad S_{\circ}^{2}-\pi_{\varpi}^{2}={\mathcal{S}}_{\perp}\,,$ (128)
and [^22] Notice that we here adopt the right-handed representation of the Marck tetrad, instead of the left-handed one widely used in the literature [cf. Refs. [65, 35, 31, 42]]; the details are explained in Appendix. C.1.
where we used $\pi_{\alpha}=p_{\alpha}+O(\epsilon)$ and evaluated the right-hand side of Eq. (127) with $S_{\theta}^{\text{\tiny{TD}}}$ and $S_{\phi}^{\text{\tiny{TD}}}$ in Marck’s representation. Here, the constant-angle $\phi_{s}$ defines the orientation of the TD spin (in the Marck frame), and ${\psi_{p}}={\psi_{p}}({\bar{\tau}})$ describes the Marck’s precession phase that grows with the proper time $\bar{\tau}$ according to the differential equation in Eq. (180).
Equation (128) immediately allows us to explain why the constant of motion $\pi_{\varpi}$ in Eq. (116) is interpreted as the “parallel component” of the spin; this is equivalent to the amplitude of the third leg of Marck tetrad [cf. Eq. (178)], aligned with the direction of the Killing-Yano angular momentum vector ${\cal L}^{a}$. Similarly, the overall amplitude $\sqrt{S_{\circ}^{2}-\pi_{\varpi}^{2}}$ in Eq. (126) can be interpreted as the “perpendicular” component of the TD spin, in the sense that it is perpendicular to ${\cal L}^{a}$.
When we insert Eq. (128) within Eq. (126), $S_{t}^{\text{\tiny{TD}}}(\tilde{x}^{i},\tilde{\pi}_{i})$ and $S_{r}^{\text{\tiny{TD}}}(\tilde{x}^{i},\tilde{\pi}_{i})$ can be directly compared with each corresponding component in Eq. (129), by covariance of the TD spin $S^{a}_{\text{\tiny{TD}}}$ (in contrast to the non-covariant $3$-spin vectors $\vec{S}$). Upon using $p_{t}=\pi_{t}+O(\epsilon),\,p_{r}=\tilde{\pi}_{r}+O(\epsilon)$ and $L=\pi_{\ell}$, this comparison leads two linear relations put in the matrix form
[…]
It is in this sense that the variable ${\varpi}$ is related to the spin precession; when the canonical variable ${\varpi}$ is rotated by the angle $\Theta$ (in the phase space $\mathcal{P}$), it coincides with the Marck precession phase $\psi_{p}$ with the phase offset $\phi_{s}$ (in the Schwarzschild spacetime).
As a conclusion of this subsection, we need to comment on the angle $\Theta$ that appears in Eq. (132). The physical meaning of this angle is currently unknown. We have attempted to derive Eqs. (131) and  (133) within our Hamiltonian formalism (i.e. without relying on the comparison with Marck’s representation of $S_{a}^{\text{\tiny{TD}}}$), but, for reasons that remain unclear to us, this approach does not seem to not seem to be effective in this case.232323 Technically, it was challenging to infer the specific combination $r^{2}\mu^{2}+\pi_{\ell}$ that enters the denominator of Eq. (131) (as well as Eq. (180)) from the canonical Hamiltonian system. The question of why the Marck’s precession phase $\psi_{p}$ must be a superposition of the canonical variable ${\varpi}$ and ”the” angle $\Theta$ (when viewed from the phase space $\mathcal{P}$) is a profound one; we leave it open for the time being.
@@ IV.3 Action-angle variables
The Liouville-Arnold theorem guarantees the existence of action-angle variables $(\vartheta^{i},\mathcal{J}_{i})$, yet the difficulty remains of determining their explicit functional forms. Once again, our previously established canonical Hamiltonian system in Sec.  IV.1 provides a practical method for achieving this objective. The key point is that four (out of five) canonical momenta $(\pi_{t},\pi_{\nu},\pi_{\ell},\pi_{\varpi})$ in Eq. (112) are all constants of motion, and the Hamiltonian $H_{c}$ in Eq. (114) is already reduced to the “radial” Hamiltonian for $(r,\,\tilde{\pi}_{r})$ essentially. The overall structure of the Hamiltonian system is therefore identical to the well-known one for the geodesic dynamics in Schwarzschild [cf. Appendix. B], except for the addition of the spin canonical pair $({\varpi},\,\pi_{\varpi})$. We can immediately obtain the action-angle variables for the spinning-body in much the same way as those defined in the case of the non-spinning body [cf. the Kerr geodesic results in Refs [46, 53]].

### v3

@@ II.3.3 Summary
where the notation $y_{\mathcal{N}}^{\textrm{A}}$ includes an A for “Andoyer”. They are in a 1-to-1 correspondence with the old canonical set $y_{\mathcal{N}}$ defined in Sec. I.2.3 via Eqs. (61) and (62), and they generalize the classical Schwarzschild geodesic elements [cf. Sec. II.1].
@@ II.4.1 Andoyer Hamiltonian
Despite the striking similarity with the first integrals of the geodesic case [cf. the discussion below Eq. (53)], we would like to emphasize two points. First, all these first integrals include a spin contribution, and are not equal to their geodesic counterpart (despite their name). Second, the constant plane, orthogonal to $\vec{L}$, is not the orbital plane, in contrast to the geodesic case. Generic spinning orbits are not planar, but they do stay $O(\epsilon)$ to the plane orthogonal to $\vec{L}$, which can be considered to be the “mean” orbital plane [cf. Fig. 3].
@@ II.4.2 Constraints and equations of motion on $\mathcal{N}$
where we made the dependence on the Rüdiger invariant $\mathfrak{K}$ [cf. Eq. (45a)] explicit, even though, on $\mathcal{N}$, this quantity is merely function of the Andoyer variables; it is neither a coordinate nor a first integral. However, it reveals the practical advantage of choosing $C^{0}$ and $C^{1}$ over other choices for the TD SSC constraints: $C^{0}$ and $C^{1}$ can be directly related with the Rüdiger invariant $\mathfrak{K}$, and will be very useful on $\mathcal{P}$.
@@ II.5.2 Hill equation for the spin
It is worth noting that the Hill equation in Eq. (71) in the equatorial limit $\theta\to\pi/2$ bears resemblance to the other Hill equation for the TD spin $S^{a}_{\text{\tiny{TD}}}$ derived by Apostolatos in Ref. [134]. This latter equation is given by [cf. Eqs. (24) and (25) in Ref. [134]]:
$\frac{\mathrm{d}^{2}(rS^{\phi}_{\text{\tiny{TD}}})}{\mathrm{d}\ell^{2}}+\left(1-\frac{3M}{r}\right)\,(rS^{\phi}_{\text{\tiny{TD}}})=0\quad\text{and}\quad-S^{r}_{\text{\tiny{TD}}}=\frac{\mathrm{d}(rS^{\phi}_{\text{\tiny{TD}}})}{\mathrm{d}\ell}\,,$ (72)
where we used $\phi=\ell$ in the equatorial limit with $\nu=0$. Indeed, Eqs. (71) and (73) are mutually consistent. In the equatorial limit, $S^{r}_{\text{\tiny{TD}}}$ and $S^{\phi}_{\text{\tiny{TD}}}$ in terms of Andoyer variables $y_{\mathcal{N}}^{\textrm{A}}$ may be manipulated to take schematic form given by
$S^{A}_{\text{\tiny{TD}}}=(\ldots)\tan s+(\ldots)\pi_{s}\quad\text{with}\quad A={(r,\phi)}\,,$ (73)
where the coefficients ()($\ldots$)( … ) are functions of $r$, $\pi_{r}$ and integrals of motion $\mu,E$ and $L$ [cf. Sec. II.4.1] Inserting this relation into Eq. (73), it then becomes evident that the result recovers Eq. (71).
Nevertheless, there is an essential distinction between the two Hill equations. Despite their formal identity, it can be demonstrated that $rS^{\phi}_{\text{\tiny{TD}}}\neq\pi_{s}$, and similarly, $S^{r}_{\text{\tiny{TD}}}$ is not just proportional to $\tan s$. In other words, the Andoyer variables $(\tan s,\pi_{s})$ are not merely components of the TD spin vector; rather, they are a non trivial combination of the latter given by Eq. (73) because the coefficients are not identically zero even in the equatorial limit. From this perspective, our Hill equation in Eq. (71) may be considered an extension of Apostolatos’s formulation to a non-equatorial orbit (and symplectic setups).
@@ III.1 Hamiltonian constraints
Two points must be made regarding the right-hand side of the aforementioned formula in Eq. (75). First, the non-vanishing 171717 The formula in Eq. (75) is singular when $\{C^{0},C^{1}\}=0$, which occurs when $\pi_{\alpha}=0$ [cf. Eq. (82)]. The singularity is only apparent, however, because both $\{C^{0},C^{1}\}$ and the numerator of the formula is always proportional to $\pi_{\alpha}$, which we will see from Eq. (190). The limit $\pi_{\alpha}\rightarrow 0$ of Eq. (75) is therefore regular. of the denominator $\{C_{0},C_{1}\}$ is a consequence of the symplectic nature of the Poisson structure on $\mathcal{P}$, as explained in Sec. V.A.2 of Paper I. Second, the right-hand side must be first computed with the brackets on $\mathcal{N}$, and then simplified the result, making use of the two constraints $C_{0}=0$ and $C_{1}=0$ if necessary. It is important to appreciate the order of these (non-commuting) operations; the redundant variables in Eq. (75) must be removed only after the brackets on $\mathcal{N}$ are completely evaluated.
@@ III.1.1 Andoyer spin variables
$\Delta^{2}$ $:=\sqrt{1-S_{\circ}^{2}/\pi_{\zeta}^{2}}\,\bigl{(}{\pi_{s}}\sin\zeta\cos s+{\pi_{\zeta}}\cos\zeta\sin s\bigr{)}\,,$ (76e)
@@
We note that these Andoyer spin variables $(\vec{\Sigma},\vec{\Delta})$ are identical to $(\vec{S},\vec{D})$ in Eq. (29), except that the pair $(\sigma,\pi_{\sigma})$ is replaced with $(s,\pi_{s})$. Since $s$ is a combination [cf. Eq. (61)] of the orbital angles $\theta,\phi$ and the spin angle $\sigma$ adapted to Schwarzschild’s spherical symmetry, $(\vec{\Sigma},\vec{\Delta})$ will be much more practical to use than $(\vec{S},\vec{D})$, as we shall see.
@@ III.1.2 Choice of 10 variables on $\mathcal{P}$
Our next step is to chose $10$ variables to cover the phase space $\mathcal{P}$. Any subset of the $12$ Andoyer variables $y_{\mathcal{N}}^{\text{A}}$ (or any combination thereof) on $\mathcal{N}$ is in principle possible, altough in practice there is no general rule as to which variables should be retained or discarded. Our approach will be to retain the constants of motion $(E,\,L_{z},\,L^{2}=\mathfrak{Q},\,\mathfrak{K})$ (or, again, any combination thereof) introduced in Sec. I.3. Indeed, since we know that the system is integrable on $\mathcal{P}$ [cf. Sec. VI in Paper I], these constants of motion can be readily promoted to momenta on $\mathcal{P}$. We just have to express them in terms of Andoyer variables. These expressions can readily be found below Eq. (67) for $(E,L_{z},L)$, while that for $\mathfrak{K}$ is simply an upgrade from Eq. (45a) to
@@ III.1.4 $\mathcal{P}$-brackets
We can now compute the $\mathcal{P}$-brackets of all pairs of the $10$ variables of $y_{\mathcal{P}}^{\text{C}}$ in Eq. (80) via Eq. (75). To compute the $\mathcal{N}$-brackets in the right-hand side in the formula, we follow the general strategy adopted in the calculation of $\{C^{0},C^{1}\}$ in the preceding section; first, apply the Leibniz rule to relate the expression to the brackets in terms of $(\Sigma^{I},\Delta^{I})$, discarding any terms that are quadratic or higher in spins. Next, employ the SO$(1,3),\mathcal{N}$-brackets in Eq. (78) to eliminate the brackets in the expression. Finally, replace $(\Sigma^{I},\Delta^{I})$ with $\Gamma$ and $\pi_{\varpi}$ (using the relation in Eq. (188) of Appendix C.1), to put the result in the form solely expressed in terms of $y_{\mathcal{P}}^{\text{C}}$ in Eq. (80). The key technical point here is to fully make use of the Andyer spin variables $(\Sigma^{I},\Delta^{I})$ in the intermediate stage, which we find to be the most optimal. [^18] In principle, it is possible to directly calculate the $\mathcal{P}$-brackets from the Andoyer variables in Eq. (66), without relying on $(\Sigma^{I},\Delta^{I})$. However, it appears that such a strategy is computationally intractable in general (even with a computer algebra system).
An explicit listing of the $\mathcal{N}$-brackets between the constraints $C^{0}$ and $C^{1}$ in Eq. (81) and the $10$ variables $y_{\mathcal{P}}^{\text{C}}$ in Eq. (80) is given in Appendix C.1. Collecting the results from Eq. (82) and Eq. (190), and then inserting them into Eq. (75), we arrive at
@@ III.1.5 Generalized law of motion
where we used the Leibniz rule at the second equality. It is then straightforward to explicitly evaluate this law of motion, inserting the $\mathcal{P}$ brackets in Eq. (88) within Eq. (90). Performing the computation, we obtain191919 As expected, we also get the trivial equations $\{\mathcal{I},H_{\mathcal{P}}\}^{\mathcal{P}}=0$ for all first integrals $\mathcal{I}\in(\nu,\pi_{t},\pi_{\ell},\pi_{\nu},\pi_{\varpi})$.
@@ III.2.4 Fixing the residual dependence
The determination of $\tilde{t}_{3},\tilde{\pi}_{3},\tilde{\ell}_{3}$ and ${\varpi}_{3}$ as a function of variables $(r,\pi_{t},\pi_{\ell},\pi_{\varpi})$ is instead based on an alternative approach, which we detail in Appendix. C.2. The main conclusion of this analysis is that there is no unique solution, due to the freedom in choosing canonical coordinates in symplectic mechanics. Nevertheless, there is a trivial homogeneous solution expressed as [cf. Eqs. (193) and (194)]
@@
$y_{\mathcal{P}}^{\text{F}}=(\tilde{t},r,\nu,\tilde{\ell},{\varpi},\pi_{t},\tilde{\pi}_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi})\,,$ (112)
@@ IV Canonical Hamiltonian system on $\mathcal{P}$
As previously stated in the introductory section .2.1, the canonical Hamiltonian system presented in Sec. IV.1 is completely equivalent to the original linear-in-spin, MPTD $+$ TD SSC system in Eqs. (12)-(16). While the canonical coordinate system $y_{\mathcal{P}}^{\text{F}}$ used in the canonical Hamiltonian system differs apparently from $(x^{\alpha},p_{\beta},S^{\gamma\delta})$ used in the original system, they are, in fact, related through a one-to-one correspondence [cf. Eq. (2)], a fact that was established through discussions in all preceding sections.
@@ IV.1 The canonical Hamiltonian system
The first five of them carry the same physical interpretations as those in the geodesic case, and only $S_{\parallel}$ appears as a new integral of motion specific to the spinning case. In this sense, $H_{c}$ is a natural generalization of the well-established (covariant) geodesic Hamiltonian introduced in, e.g., Refs. [68, 75] [cf. Eq. (156) in Appendix. A]. It is important to note that only five out of the six integrals of motion are in pairwise involution,  [^20] Two first integrals $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are in involution whenever $\{\mathcal{I}_{1},\mathcal{I}_{2}\}=0$. because the canonical pair $(\nu,\,\pi_{\nu})$ satisfies the bracket $\{\nu,\pi_{\nu}\}=1$ by definition.
@@ IV.3 Action-angle variables
The Liouville-Arnold theorem guarantees the existence of action-angle variables $(\vartheta^{i},\mathcal{J}_{i})$, yet the difficulty remains of determining their explicit functional forms. Once again, our previously established canonical Hamiltonian system in Sec.  IV.1 provides a practical method for achieving this objective. The key point is that four (out of five) canonical momenta $(\pi_{t},\pi_{\nu},\pi_{\ell},\pi_{\varpi})$ in Eq. (115) are all constants of motion, and the Hamiltonian $H_{c}$ in Eq. (117) is already reduced to the “radial” Hamiltonian for $(r,\,\tilde{\pi}_{r})$ essentially. The overall structure of the Hamiltonian system is therefore identical to the well-known one for the geodesic dynamics in Schwarzschild [cf. Appendix A], except for the addition of the spin canonical pair $({\varpi},\,\pi_{\varpi})$. We can immediately obtain the action-angle variables for the spinning-body in much the same way as those defined in the case of the non-spinning body [cf. the Kerr geodesic results in Refs [68, 75]].